Create a function named determineList which inputs two characters of dimensions 1×1 which represent the start and end of the range. These two character datatype inputs can be of either alphabetic (A-Z,a-z) or digits (0-9) type. The function should then output a list (character array of dimension N × 1). (in matlab)
Answer #1

MATLAB CODE:

addpath(pwd)
function[arr] = determineList(s,e)
N = e-s+1
arr =zeros(N,1)
for i = 1:N
arr(i) = (s+i-1)
end
arr = char(arr)
end
D = (determineList('A','C'))

OUTPUT:

Code Screenshot: